A jar on your desk contains 12 black, 8 red, 10 yellow, and 5 green jellybeans. You pick a jellybean without looking. Find the probability that it is red or yellow.
Answer by ultimate44(2)






Say;






black jellybeans=12,






red jellybeans=8,






yellow jellybeans=5,






green jellybeans=5,






total jellybeans=12+8+5+5=30






Therefore:






P(red jellybeans)=8/30.. while p(yellow jellybeans)=5/30






Hence.p(red or yellow)=(8/30)+(5/30)=13/30